# Improving Socratic Question Generation using Data Augmentation and Preference Optimization

### v1

@@ 5 Experimental Settings
In this section, we detail the implementation setup, methods compared, and metrics used to evaluate Socratic question generation.
@@
Implementation details. In the data augmentation phase, we query OpenAI’s [^1] https://openai.com/ GPT-4 using a rate-based API. We set the invalid Socratic questions data generation parameters with a temperature of 0.5 to encourage moderate randomness in the outputs, we also set the maximum number of tokens to be generated to 2000. For the consistency checking GPT-4 model, we use a temperate of 0 to maintain determinism and set the maximum tokens to be generated to 200. In the preference optimization phase, we use Code-Llama [24] pre-trained for instruction following tasks, particularly on code data [^2] https://huggingface.co/codellama/CodeLLama-7b-hf. We load our Code-Llama model in an 8-bit configuration and train using QLora [7] with the peft [^3] https://huggingface.co/docs/peft HuggingFace library. For the SFT step, we fine-tune the model for 5 epochs with a learning rate of 3e-5, and a batch size of 2 by accumulating gradients for creating a virtual batch size of 64 which takes about 10 hours to train. For the DPO step, we fine-tune the model for 1 epoch with a learning rate of 3e-5 and a $\beta$ (which denotes the KL-loss [12] between the preference policy learned and the reference SFT policy) of 0.1, with a batch size similar to that of SFT, which takes about 6 hours to train. We conduct our experiments on a single Nvidia A6000 GPU with 48G of volatile memory. For the DPO experiments, we carry out a grid search using hyperparameters learning rate as 1e-5, and 3e-5, $\beta$ of 0.1, and 0.5 and number of epochs as 1 and 2 to arrive at the best-performing hyperparameters as mentioned above.
@@
Methods. To decode the trained LLM output, we use two decoding techniques, greedy and nucleus sampling, with a $p$ value of 0.9 temperature of 1, and the number of return sequences of 5 and 10. We refer to these methods coupled with the trained SFT method as SFT Greedy, SFT Sample-5, SFT Sample-10, and similarly for the DPO methods. Greedy decoding takes 30 minutes to complete, whereas Sample-5 takes an hour and Sample-10 decoding takes about 2 hours.
@@
Metrics. To measure the similarity between the generated Socratic questions and the ground truth questions, we use three commonly used evaluation metrics in natural language generation tasks: BLEU-4 [19], which measures the n-gram overlap between the generated and ground-truth questions, BERTScore [36] based on the DeBERTa language model [10], which measures the semantic similarity, and Rouge-L [17], which measures n-gram overlap based on the longest common subsequence (LCS). In addition, the dataset we use [1, 2] provides multiple ground truth Socratic questions at each dialogue turn. To measure the similarity between a set of $n$ LLM-generated questions with a set of $m$ ground truth questions, we adopt the process used in [1], which uses Edmond Blossom algorithm [9] to find the maximum matching in a complete bipartite graph between the two sets with a total of $mn$ edges, where the weight of each edge is computed using one of the metrics mentioned above. This step guarantees that every ground-truth question corresponds to, at most, one LLM-generated question, inhibiting semantically equivalent LLM generations from artificially inflating the metric scores.
@@ 6 Results and Discussions
Table 1: Performance comparison across different methods. All GPT baseline results are reported in [1]. Best results are bolded (for R), underlined (for F1), and colored.
Method Rouge-L BLEU-4 BERTScore
P R F1 P R F1 P R F1
GPT-3.5 21 14.3 17 3.2 1.7 2.0 56.0 43.5 48.9
GPT-3.5 + CoT 20.3 9.7 12 2.3 0.8 1.1 61.7 35.8 41.6
GPT-4 14.1 23.3 17.6 3.2 4.3 3.6 35.4 62.6 45.2
GPT-4 + CoT 5.2 26.6 8.1 0.9 4.8 1.4 12.6 64.8 19.5
SFT Greedy 29.66 13.41 17.19 2.0 0.91 1.18 61.77 29.32 36.77
DPO Greedy 30.56 13.26 17.12 2.5 1.06 1.38 65.86 32.69 40.28
SFT Sample-5 14.11 25.95 17.11 1.10 2.08 1.36 32.09 62.85 41.06
DPO Sample-5 15.12 27.93 18.31 1.19 2.28 1.46 34.77 64.26 42.00
SFT Sample-10 9.01 32.28 13.36 0.719 2.61 1.08 19.02 67.45 27.95
DPO Sample-10 9.36 34.49 13.91 0.787 3.0 1.18 19.1 68.45 28.11
Table 1 shows the comparison between different methods on the metrics defined for our task. All the GPT-3.5 and GPT-4 results are taken from what was reported in prior work [1]. The most important metrics here are the recall scores (R), since [1] carried out a human evaluation to show that they are positively correlated with human ratings while precision (P) and F1 are not. We see that DPO Sample-10 has the highest recall score on the Rouge-L and BERTScore metrics, higher by 7.89 and 3.65 points on Rouge-L and BERTScore, respectively, compared to GPT-4 + CoT, which is the best among GPT-based methods. We note that although the recall of our methods is higher, their F1 scores are lower than some of the GPT-based methods. The reason is that finding the maximum bipartite matching (as detailed above) penalizes LLM over-generation by classifying them as false positives and hence decreasing the precision. Plus, we see that GPT + CoT has the highest recall scores and the lowest F1 score among all the GPT-based methods, which is consistent with results in [1] showing that humans prefer the GPT-4 + CoT outputs. From this analysis, we can conclude that by obtaining higher recall scores, our methods outperform existing state-of-the-art prompt-based methods.
We also see that our methods, despite using an open-source model with 7B parameters, are better than or comparable to propriety models like GPT-3.5-turbo and GPT-4, which have many more parameters. This result is important since it means that our method can be much more scalable and cost-effective than methods relying on GPT models, without sacrificing any performance degradation, making it available to students who do not have access to proprietary models. We note that the BLEU scores of the SFT and DPO methods are lower than the GPT-based methods despite the Rouge-L and BERTScore results being higher (or comparable). The reason is that although the questions generated by SFT and DPO methods are semantically similar to the ground truth, they do not have the exact lexical matches as seen in the ground-truth questions, while BLEU focuses more on the precision of the generated questions.
We also see that DPO consistently outperforms SFT across all metrics, which justifies the need to perform preference optimization using a preference dataset obtained through data augmentation. The gap in the results between DPO and SFT narrows down as we increase the number of generated questions, i.e., in terms of the difference in the BERTScore recall, the gap is the highest for Greedy (3.37), followed by Sample-5 (1.41), and Sample-10 (1.0). The reason is that as the number of generated questions increases, the possibility that some of them are similar to the ground truth questions also increases.
@@
Problem
Write a function “top_k(lst: List[int], k: int) -> List[int]” that returns the top k largest elements in the list. You can assume that k is always smaller than the length of the list.
@@
Buggy Code
@@
Bug Description
The function removes the element at index ‘max(lst)’ instead of removing an element equal to ‘max(lst)’. Consequently, the function throws an IndexError on line 5 when a removed value in ‘lst’ is greater than the length of ‘lst’.
Bug Fixes
On line 5, replace ‘lst.pop(max(lst))’ with ‘lst.remove(max(lst))’
Conversation
Student: Hi. I am confused. My code doesn’t seem to work. Can you help?
Ground Truth
1. Ok, no worries. Let’s review your code line by line. Could you please explain it to me?
3. Could you please explain what line 5 in your code does?
Generated
Irrelevant: What happens if you enter an empty list as the input?
Questions
Repeated: Do you know what might be the issue?
7 Case Study
We now use a case study to illustrate why our method leads to better Socratic question generation. Table 2 shows an example of the augmented data, i.e., invalid questions generated by GPT-4 for an example problem, which asks students to write code to return the largest k elements in a list. The student’s code incorrectly removes elements at index max(lst) as opposed to removing elements equal to max(lst), thereby causing an IndexError. The potential fix to the code is to replace the .pop() function with .remove(). In the conversation, we see that the student knows the problem lies in their use of .pop(). The ground truth Socratic questions for this dialogue turn are highly generic, asking the student to review the code line by line, apply an example test case, or do further reading on Python documentation. We see that the four types of invalid questions generated by GPT-4 are: the irrelevant question is out of context and does not help the student understand the bug in their code. The repeated question has already been mentioned by the instructor. The direct questions reveal the problematic function .pop() and do not give room for the students to discover the problem themselves. The premature question directly suggests a code change to replace the .pop() with .remove() function even before the student has realized the actual bug. These diverse examples of invalid questions serve as good training data to let an LLM know what kinds of invalid questions it should avoid generating.
Table 3 shows the questions generated by different methods for an example problem that asks students to write code that returns the factorial of a number. The student’s bug has some indexing errors with the range function, resulting in the fact variable being multiplied by 0 when the loop starts, hence resulting in the output of 0 no matter what the input is. In the conversation, we see that the ground truth questions encourage the student to debug the code by printing the value of lines 5 and 6 to examine the variables fact and i along with the role of range(n). The SFT Greedy output correctly asks the student the value of i but does not provide more details, such as the value of n. The DPO Greedy method is more helpful since it asks the student to check the value of fact specifically for n=1. The first and fifth questions by SFT Sample-5 are invalid and directly ask the value of fact when i is 0, which directly reveals the problem that fact is always 0. The third and fourth outputs are either irrelevant or repeated. The second question, which asks the student to examine the value of fact is valid since it does not directly reveal the bug. In contrast, most of the DPO Sample-5 questions are valid, since they urge the student to examine the value of i and fact on lines 5 and 6 with specific values of n, without directly revealing the bug that i is always 0. Through these two comparisons, we see that DPO improves Socratic question generation compared to SFT and that DPO Sample-5 is highly capable of generating valid yet diverse questions.
Table 3: Socratic questions generated by different methods.
Problem
Write a function ‘factorial(n:int) -> int’ that computes the factorial n! of a natural number n, which is defined mathematically as: $0!=1$ and $n!=n\times(n-1)!$. Additionally, if the input integer n is negative the function should return 0.
Buggy Code
@@ 7 Case Study
Bug Description
On line 6, ‘fact’ is multiplied with 0 in the first iteration of the for loop. Consequently, at every iteration fact stays equal with 0 instead of being updated to be equal with the factorial of ‘(i + 1)’. Therefore, the function will return 0, irrespective of n
Bug Fixes
Replace ‘i’ with ‘(i + 1)’ in line 6. Replace ‘range(n)’ with ‘range(1, n + 1)’ in line 5.
Conversation
Student: Hi! I implemented the factorial function but it doesn’t work and I do not know why. Can you help?
Student: For n = 1 or larger it returns the same value, 0.
Ground Truth
1. Let’s see what happens when ‘n’ is 1. What is the first value that is assigned to variable ‘i’ in line 5?
4. Let’s see what happens when ‘n’ is 1. What does ‘range(n)’ do when ‘n’ is 1?
5. Can you tell me what ‘range(n)’ does?
SFT Greedy
What is the value of ‘i’ in the first iteration of the for loop?
DPO Greedy
Okay. Let’s start with ‘n = 1’. What do you think ‘fact’ is on line 6?
SFT Sample-5
1. What is the value of ‘fact’ at line 6 when ‘i’ is 0?
4. What does this tell us about the implementation?
5. Are you multiplying the initial ‘fact’ value of 1 by ‘i’ at each step?
DPO Sample-5
1. Alright, so for any value of ‘n’ that we pass to the function, what is the output of line 6?
2. Can you tell me what value of ‘i’ is returned in line 6 when ‘n = 1’?
4. Okay, let’s go line by line. Let’s start from line 5. What do we see?
@@ 8 Conclusions and Future Work
In this work, we proposed a method for Socratic question generation in programming problem feedback scenarios. Our method consists of a data augmentation phase to create a preference dataset by synthetically generating invalid questions according to four possible ways. We then use this preference dataset to fine-tune an open-source LLM, Llama 2, with 7B parameters, using direct preference optimization (DPO) using a Socratic questioning dataset that consists of student-instructor conversations in code debugging scenarios. Our results show that the preference-optimized 7B LLama 2 model significantly outperforms existing state-of-the-art prompting methods (on common text similarity metrics) that rely on much larger GPT models (25 times larger), by avoiding invalid questions after training on the augmented dataset. Our method paves the way toward an open-source, accessible, cheaper, privacy-preserving, and yet effective alternative to generating Socratic questions to improve students’ learning outcomes by not having to rely on proprietary rate-based API-accessible models like GPT-4.
There are several avenues for future work. First, we can increase the precision of our method by overgenerating Socratic questions and then rank them to select the top-k questions [3], to reduce the number of false positives. Second, we can develop a technique to differentiate different types of invalid Socratic questions and not treat them equally while performing preference optimization. This technique would require us to modify the inherent objective function of DPO to incorporate more than one unpreferred question for a single preferred question, which may give us fine-grained control over the LLM generations. Third, we can experiment with open-source LLMs that are larger than 7B to see whether DPO provides more significant gains over SFT on larger models on the Socratic question generation task. Finally, we can experiment with alternative preference optimization methods, such as KTO [8] which do not need explicit preference data in the form of pairs of valid and invalid questions.

### v3

@@ 5 Experimental Settings
In this section, we detail the implementation setup, methods compared, and metrics used to evaluate our Socratic question generation method.
@@
Implementation details. In the data augmentation phase, we query OpenAI’s [^2] https://openai.com/ GPT-4 using a rate-based API. We set the temperature of the GPT-4 model to 0.5 to encourage moderate randomness in the outputs. For the consistency checking GPT-4 model, we use a temperate of 0 to maintain determinism. In the preference optimization phase, we use Code-Llama (7B) Roziere et al. (2023) pre-trained for instruction following tasks, particularly on code data [^3] https://huggingface.co/codellama/CodeLLama-7b-hf. We load our Code-Llama model in an 8-bit configuration and train using QLora Dettmers et al. (2023) with the peft [^4] https://huggingface.co/docs/peft HuggingFace library to facilitate efficient fine-tuning. For the SFT step, we fine-tune the model for 5 epochs with a learning rate of 3e-5, and a batch size of 2 by accumulating gradients for creating a virtual batch size of 64 which takes about 10 hours to train on a single Nvidia A6000 GPU. For the DPO step, we fine-tune the model for 1 epoch with a learning rate of 3e-5 and a $\beta$ (which denotes the KL-loss Joyce (2011) between the preference policy learned and the reference SFT policy) of 0.1, with a batch size of 2, which takes about 6 hours to train. For the DPO experiments, we carry out a grid search using hyperparameters learning rate as 1e-5, and 3e-5, $\beta$ of 0.1, and 0.5 and number of epochs as 1 and 2 to arrive at the best-performing hyperparameters as mentioned above.
Methods. As baselines, we perform zero-shot prompting of the LLama 2 Chat model [^5] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf Touvron et al. (2023), denoted by LLama, to generate all possible Socratic questions for the current conversation turn. We also prompt LLama 2 in a chain-of-thought Wei et al. (2022) manner to first generate the current student misconceptions and then generate the Socratic questions, denoted by LLama (CoT).
@@
To decode our trained (SFT and DPO) LLM, we use two decoding techniques, greedy and nucleus sampling, with a $p$ value of 0.9 temperature of 1, and a number of return sequences of 5. We refer to these methods coupled with the trained SFT method as SFT Greedy, SFT Sample-5, and similarly for the DPO methods. Greedy decoding takes 30 minutes to complete, whereas Sample-5 takes an hour.
@@
Metrics. To measure the similarity between the generated Socratic questions and the ground truth questions, we use two commonly used evaluation metrics in natural language generation tasks: BERTScore Zhang* et al. (2020) based on the DeBERTa language model He et al. (2021), which measures the semantic similarity, and Rouge-L Lin (2004), which measures n-gram overlap based on the longest common subsequence (LCS). In addition, the dataset we use Al-Hossami et al. (2024, 2023) provides multiple ground truth Socratic questions at each dialogue turn. To measure the similarity between a set of $m$ LLM-generated questions with a set of $n$ ground truth questions, we adopt the process used in Al-Hossami et al. (2024), which uses Edmond Blossom algorithm Galil (1986) to find the maximum matching in a complete bipartite graph between the two sets with a total of $mn$ edges, where the weight of each edge is computed using one of the metrics mentioned above. This step guarantees that every ground-truth question corresponds to, at most, one LLM-generated question, inhibiting semantically equivalent LLM generations from artificially inflating the metric scores. The number of True Positives (TP) is the total sum of the weights of all edges in the optimal matching. False Positives (FP) are calculated by summing the difference between every weight of an edge in the matching with 1. Any unmatched LLM-generated question counts 1 towards False Positive. Similarly, any unmatched ground truth question counts 1 towards False Negative (FN). The TP, FP, and FN values are used to compute the precision, recall, and F1 score for a particular metric. The metric penalizes over-generated LLM questions that do not match with any ground truth questions by classifying them as an FP, thus decreasing the precision.
@@ 6 Results and Discussions
Table 1: Performance comparison across different methods. All GPT baseline results are reported in Al-Hossami et al. (2024). Boldface represents the highest value/s for that column.
Method Rouge-L BERTScore
P R F1 P R F1
GPT-3.5 21.0 14.3 17.0 56.0 43.5 48.9
GPT-3.5 (CoT) 20.3 9.7 12.0 61.7 35.8 41.6
GPT-4 14.1 23.3 17.6 35.4 62.6 45.2
GPT-4 (CoT) 5.2 26.6 8.1 12.6 64.8 19.5
LLama 12.8 18.6 13.2 36.0 48.3 35.9
LLama (CoT) 13.7 15.5 13.2 42.3 49.0 41.0
SFT Greedy 29.7 13.4 17.2 61.8 29.3 36.8
DPO Greedy 30.6 13.3 17.1 65.9 32.7 40.3
SFT Sample-5 14.1 26.0 17.1 32.1 62.9 41.1
DPO Sample-5 15.1 27.9 18.3 34.8 64.3 42.0
In the consistency checking step of the data augmentation phase, we see that 72% of the generated questions are considered for the preference dataset creation as 27% of the generated questions are classified as “good” and 1% as “incorrect”. This result shows that GPT-4 is more prone to generate “good” questions for particular dialogue turns than incorrect questions that do not relate to the problem and the dialogue.
For the task of Socratic question generation, Table 1 shows the comparison between different methods on the metrics defined for our task. All the GPT-3.5 and GPT-4 results are taken from prior work Al-Hossami et al. (2024). We observe that GPT-4 (CoT) has the highest recall and yet the lowest F1 score. This observation is because, GPT-4 generates a large number of Socratic questions a few of which are similar to the ground truth questions, however, a significant fraction of the generated questions do not correspond to any ground truth questions, hence being labeled as false positive, thus decreasing the precision. Al-Hossami et al. (2024) also carry out manual analysis to show that GPT (CoT) outputs are the best despite having low F1 scores. This observation can be attributed to the fact that GPT (CoT) has the highest recall among all other GPT methods and hence better corresponds to the ground truth questions.
For the baseline methods that use zero-shot LLama prompting, we observe that LLama (CoT) is the best, which shows that chain-of-thought prompting to first generate the students’ current misconceptions followed by the Socratic questions is effective. Among the preference optimization experiments, we see that DPO consistently outperforms SFT. We also observe that the LLama (CoT) performs as well as DPO Greedy in terms of BERTScore F1 as LLama (CoT) generates a higher number of Socratic questions whereas the DPO Greedy method just generates one. Hence, the recall of the DPO Greedy method is lower than that of LLama (CoT). Among decoding variants, we see that the Sample-5 method is better than the Greedy method highlighting the importance of sampling multiple possible Socratic questions instead of just one.
Overall, we see that our preference-optimized models with DPO give the best Rouge-L scores for all precision, recall, and F1 scores with DPO Greedy having the highest precision and DPO Sample-5 having the highest recall and F1 score among all the methods. DPO Greedy has the highest BERTScore precision, whereas DPO Sample-5 has a recall comparable to the best GPT method, GPT-4 (CoT). These results suggest that the DPO-optimized LLama 2-7B model is better than (or as effective as) much larger models like GPT-4 (25 times larger) for Socratic question generation.
@@
Problem Write a function “top_k(lst: List[int], k: int) -> List[int]” that returns the top k largest elements in the list. You can assume that k is always smaller than the length of the list.
@@
Bug Description The function removes the element at index ‘max(lst)’ instead of removing an element equal to ‘max(lst)’. Consequently, the function throws an IndexError on line 5 when a removed value in ‘lst’ is greater than the length of ‘lst’.
Bug Fixes On line 5, replace ‘lst.pop(max(lst))’ with ‘lst.remove(max(lst))’
Conversation Student: Hi. I am confused. My code doesn’t seem to work. Can you help?
Ground Truth 1. Ok, no worries. Let’s review your code line by line. Could you please explain it to me?
3. Could you please explain what line 5 in your code does?
Invalid Generated Irrelevant: What happens if you enter an empty list as the input?
Questions Repeated: Do you know what might be the issue?
Table 3: Socratic questions generated by different methods.
Problem Write a function ‘factorial(n:int) -> int’ that computes the factorial n! of a natural number n, which is defined mathematically as: $0!=1$ and $n!=n\times(n-1)!$. Additionally, if the input integer n is negative the function should return 0.
Bug Description and Fix On line 6, ‘fact’ is first multiplied with 0 in the first iteration of the for loop and hence remains 0 throughout the loop. To fix the bug, replace ‘i’ with ‘(i + 1)’ in line 6 and replace ‘range(n)’ with ‘range(1, n + 1)’ in line 5.
Conversation Student: Hi! I implemented the factorial function but it doesn’t work and I do not know why. Can you help?
Student: For n = 1 or larger it returns the same value, 0.
Ground Truth 1. Let’s see what happens when ‘n’ is 1. What is the first value that is assigned to variable ‘i’ in line 5?
LLama So, it seems that the issue is that the value of ‘fact’ is not being updated correctly in the loop. Can you tell me why do you think ‘fact‘ is not being updated?
LLama (CoT) Can you explain why you think the bug is in that specific line?
SFT Greedy What is the value of ‘i’ in the first iteration of the for loop?
DPO Greedy Okay. Let’s start with ‘n = 1’. What do you think ‘fact’ is on line 6?
SFT Sample-5 1. What is the value of ‘fact’ at line 6 when ‘i’ is 0?
4. What does this tell us about the implementation?
5. Are you multiplying the initial ‘fact’ value of 1 by ‘i’ at each step?
DPO Sample-5 1. Alright, so for any value of ‘n’ that we pass to the function, what is the output of line 6?
2. Can you tell me what value of ‘i’ is returned in line 6 when ‘n = 1’?
4. Okay, let’s go line by line. Let’s start from line 5. What do we see?
Table 4: Buggy codes corresponding to the case studies
Code 1:
@@ 2. Can you tell me what value of ‘i’ is returned in line 6 when ‘n = 1’?
Code 2:
@@
7 Case Study
We now use a case study to illustrate why our method leads to better Socratic question generation. First, we show an example of invalid Socratic questions generated by our data augmentation phase. Second, we compare different methods for Socratic question generation.
Table 2 shows an example of the augmented data, i.e., invalid questions generated by GPT-4 for an example problem, which asks students to write code to return the largest k elements in a list. The student’s code (Table 4 Code 1) incorrectly removes elements at index max(lst) as opposed to removing elements equal to max(lst), thereby causing an IndexError. The potential fix to the code is to replace the .pop() function with .remove(). In the conversation, we see that the student knows the problem lies in their use of .pop(). The ground truth Socratic questions for this dialogue turn are highly generic, asking the student to review the code line by line, apply an example test case, or do further reading on Python documentation. We see that the four types of invalid questions generated by GPT-4 are: the irrelevant question is out of context and does not help the student understand the bug in their code. The repeated question has already been mentioned by the instructor. The direct questions reveal the problematic function .pop() and do not give room for the students to discover the problem themselves. The premature question directly suggests a code change to replace the .pop() with .remove() function even before the student has realized the actual bug. These diverse examples of invalid questions serve as good training data to let an LLM know what kinds of invalid questions it should avoid generating.
Table 3 shows the questions generated by different methods for an example problem that asks students to write code that returns the factorial of a number. The student’s code (Table 4 Code 2) has some indexing errors with the range function, resulting in the fact variable being multiplied by 0 when the loop starts, hence resulting in the output of 0 no matter what the input is. In the conversation, we see that the ground truth questions encourage the student to debug the code by printing the value of lines 5 and 6 to examine the variables fact and i along with the role of range(n). The LLama output is very verbose and directly reveals that the problem is in the updation of the ‘fact’ variable. The LLama (CoT) output is very vague as it does not refer to the exact line of the code. The SFT Greedy output correctly asks the student the value of i but does not provide more details, such as the value of n. The DPO Greedy method is more helpful since it asks the student to check the value of fact specifically for n=1. The first and fifth questions by SFT Sample-5 are invalid and directly ask the value of fact when i is 0, which directly reveals the problem that fact is always 0. The third and fourth outputs are either irrelevant or repeated. The second question, which asks the student to examine the value of fact is valid since it does not directly reveal the bug. In contrast, most of the DPO Sample-5 questions are valid, since they urge the student to examine the value of i and fact on lines 5 and 6 with specific values of n, without directly revealing the bug that i is always 0. Through these comparisons, we see that DPO improves Socratic question generation compared to SFT and that DPO Sample-5 is highly capable of generating valid yet diverse questions.
@@ 8 Conclusions and Future Work
In this work, we propose a method for Socratic question generation in programming problem feedback scenarios. Our method consists of a data augmentation phase to create a preference dataset by synthetically generating invalid questions according to four possible categories. We then use this preference dataset to fine-tune an open-source LLM, LLama 2-7B, using direct preference optimization (DPO). Our results show that the preference-optimized LLama 2-7B model often outperforms existing state-of-the-art prompting methods (on common text similarity metrics) that rely on much larger GPT models (25 times larger), by avoiding invalid questions after training on the augmented dataset. Our method paves the way toward an open-source, accessible, cheaper, privacy-preserving, yet effective alternative to generating Socratic questions which can improve students’ learning outcomes without having to rely on proprietary rate-based API-accessed models like GPT-4. There are several avenues for future work. First, we can develop a technique to differentiate types of invalid Socratic questions and not treat them equally while performing preference optimization. This technique would require us to modify the inherent objective function of DPO to incorporate more than one unpreferred question for a single preferred question, which may give us fine-grained control over the LLM generations. Second, we can experiment with open-source LLMs that are larger than 7B to see whether DPO provides more significant gains over SFT on larger models on the Socratic question generation task. Third, we can perform a systematic human evaluation to compare the performance of our proposed method with other baselines. Also, we can focus on designing an automatic metric (based on LLM prompting Liu et al. (2023)) other than Rouge and BERTScore which captures the helpfulness of the Socratic questions without heavily relying on assigning higher scores only to questions that have high lexical overlap with the ground-truth questions. Fourth, we can experiment with alternative preference optimization methods, such as KTO Ethayarajh et al. (2024) which do not need explicit preference data in the form of pairs of valid and invalid questions. Fifth, we can also explore if Socratic question generation helps in improving other tasks in computer science education like test case generation Kumar and Lan (2024) by posing the problem as answering several Socratic sub-questions Shridhar et al. (2022). Finally, we can also explore how to make Socratic question generation knowledge-aware, i.e., generating different questions for students with different knowledge states, which can be estimated using the open-ended knowledge tracing method for computer science education Liu et al. (2022).
9 Limitations
Our work proposes a method for preference optimizing open-source LLMs like LLama 2 for the task of Socratic question generation for student code debugging. We use only LLama 2 as the base model for carrying out preference optimization, and not other open-source models like Mistral Jiang et al. (2023). Since our main contribution is the data augmentation and preference optimization method, we use only one of the best models open-source models (LLama 2) to show that our method outperforms state-of-the-art models like GPT-4. Future work can also explore the performance of different open-source models using a variety of optimization methods including our data augmentation and preference optimization method for Socratic question generation. Also, we do not formally analyze any biases that exist in the generated augmenting data or the generated Socratic questions. Future work can focus on measuring such biases to make our methods that use these LLMs more inclusive for all students belonging to different demographics.
10 Ethics Policy
Since our invalid questions are generated using an LLM potential linguistic or cultural bias related to the pre-training of the LLM might be reflected. However, we hypothesize that this bias would be minimal as Socratic questions are goal-driven, concise, and framed in the second-person perspective directed toward the student. Our work focuses on open-source LLMs like LLama for Socratic question generation as compared to rate-based API-accessed models like GPT-4 (which is used only once during data augmentation) which implies that our methods are privacy-preserving and there is minimal chance of leakage of students’ confidential data. However, training LLMs like LLama on GPUs like A100 for 10 hours results in the emission of CO2 which might not be environmentally friendly.
11 Acknowledgments
We thank Hunter McNichols, Jaewook Lee, Alexander Scarlatos, and Nigel Fernandez for their helpful discussions around this work. The authors are partially supported by the NSF under grant DUE-2215193.
Supplementary Material
Appendix A Prompts
Table 5 shows the prompt used in the data augmentation phase for generating invalid Socratic questions. Table 6 shows the prompt used in the data augmentation phase for consistency checking to classify the generated Socratic questions into six categories. For brevity, we do not include the few-shot examples, which can be found in the code repository.
Table 5: GPT-4 Prompt for Data Augmentation - Generating Invalid Socratic questions
SYSTEM: You are a “bad” instructor (Assistant) who generates “bad” Socratic questions (based on a predefined category) to help a student debug his code in a conversation with the student. Inputs: 1. Problem Description 2. Test Cases 3. Student’s buggy code 4. Bug Description 5. Bug Fixes 6. Conversation so far Bad Categories: 1. Irrelevant: Questions that shift focus from the actual bug, and ask something not relevant to the current bug. 2. Repeated: Questions already asked or answered in the dialogue. 3. Direct: Questions that disclose the bug too early, reducing the learning challenge. 4. Premature: Questions that guide learners to code changes before they identify the issue, potentially causing confusion.
Generate “bad” Socratic questions corresponding to each of the four categories and mention the reasoning for the same.
Table 6: GPT-4 Prompt for Data Augmentation - Consistency Checking
SYSTEM: Your task is to output a probability distribution over the labels describing the category of a Socratic question generated by an assistant. Inputs: 1. Problem Description 2. Test Cases 3. Student’s buggy code 4. Bug Description 5. Bug Fixes 6. Conversation so far Bad Categories: 1. Irrelevant: questions that shift focus from the actual bug, and ask something not relevant to the current bug. 2. Repeated: questions already asked or answered in the dialogue. 3. Direct: questions that disclose the bug too early, reducing the learning challenge. 4. Premature: questions that guide learners to code changes before they identify the issue, potentially confusing. NOTE: the difference between direct and premature lies in the fact that premature questions specify code changes related to the bug whereas direct questions just reveal the bug directly. 5. Good: Questions that are subtle and naturally flow from the conversation without revealing too much about the bug directly/ suggesting code changes prematurely. 6. Incorrect: Questions that are completely out-of-context and are not related to the given problem at all.
The question whose category is to be determined is labeled as ‘Assistant Socratic Question’ in the dialogue.
Output a dictionary containing the labels as the key and the corresponding probability weights as the values. For example, Output: ‘Irrelevant’: 0.6, ‘Repeated’: 0.2, ‘Direct’: 0.1, ‘Premature’: 0.05, ‘Good’: 0.05, ‘Incorrect’:0